# Sparse-grid polynomial interpolation approximation and integration for parametric and stochastic elliptic PDEs with lognormal inputs

### v14

@@ 4 Integration
$\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}v\right\|_{X^{1}}\ \leq\ \left\|v_{\rm ev}-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}\right\|_{{\mathcal{L}}_{1}(X^{1})}\ \leq\ C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}.$
@@
$G_{2}(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq\xi\big\}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>1/q_{2}-1/2,\end{cases}$
@@ 28    $k\leftarrow k+1$
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{\rm ev}(\xi)$ be the set defined by (3.10) for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$. Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}v\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.6)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.7)
@@
Observe that the rate in (5.6) and (5.7) can be improved as $\min(\alpha,\frac{2}{q_{1}}\frac{\alpha}{\alpha+\delta})$ if the sequences $(\|u_{\boldsymbol{s}}\|_{V})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ and $(\|u_{\boldsymbol{s}}\|_{W^{r}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ have $\ell_{p_{1}}$- and $\ell_{p_{r}}$-summable majorant sequences, respectively, where $1/p_{1}=1/q_{1}+1/2$ and $1/p_{r}=1/q_{r}+1/2$. Similarly to ${\mathcal{I}}_{G(\xi_{n})}$, the quadrature operator ${\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}$ can be represented in the form of a multilevel integration method with $k_{n}$ levels:
${\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\ =\ \sum_{k=0}^{k_{n}}\delta_{k}Q_{\Lambda_{{\rm ev};k}(\xi_{n})},$
@@
$\Lambda_{{\rm ev};k}(\xi):=\ \begin{cases}\big\{{\boldsymbol{s}}\in{\mathbb{F}}_{\rm ev}:\,\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq 2^{-k}\xi\big\}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big\{{\boldsymbol{s}}\in{\mathbb{F}}_{\rm ev}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>1/q_{2}-1/2.\end{cases}$
@@
Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $\Lambda_{\rm ev}(\xi)$ be the set defined by (3.45) for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (5.8)
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.9)
@@
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{2}(\xi)$ be the set defined as in (LABEL:G_2(xi)) for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$. Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}v\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.10)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.11)
@@
Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $\Lambda_{2}(\xi)$ be the set defined by (4.11) for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (5.12)
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.13)
@@
(i) As noticed in Section 4, the sparsity of the grids $\Gamma(\Lambda_{{\rm ev};0}(\xi))$ and $\Gamma(\Lambda_{\rm ev}(\xi))$ of the evaluation points in the quadrature operators ${\mathcal{Q}}_{G_{\rm ev}(\xi)}$ and $Q_{\Lambda_{\rm ev}(\xi)}$ are much higher than the sparsity of the grids $\Gamma(\Lambda_{0}(\xi))$ and $\Gamma(\Lambda(\xi))$ of the evaluation points in the generating interpolation operators ${\mathcal{I}}_{G(\xi)}$ and $I_{\Lambda(\xi)}$.
@@
(iii) Since the use and analysis of non-adaptive construction methods for sparse-grid interpolation are important, let us compare in details our methods in Corollary 5.2 with those which has been also discussed in [4]. To construct a quadrature of the form $Q_{\Lambda}$, the author of the last work used the set $\Lambda_{m}\subset{\mathbb{F}}$ of all indices $\boldsymbol{s}$} (including non-even) corresponding to the $m$ smallest values of $\sigma_{\boldsymbol{s}}$. The number $n=n(m)$ of quadrature points in $\Lambda_{m}\subset{\mathbb{F}}$ is estimated as $n\leq Cm^{2}$ [19, Proposition 3.16]. This lead to the rate $\frac{1}{2}(1/q-1/2)$. In the present paper, we used the set $\Lambda_{\rm ev}(\xi)\subset{\mathbb{F}}_{\rm ev}$ of all only even indices $\boldsymbol{s}$} by thresholding $\sigma_{\boldsymbol{s}}\leq\xi^{1/q}$. Formally, this is similar to choosing all even indices $\boldsymbol{s}$} corresponding to the smallest values of $\sigma_{\boldsymbol{s}}$ satisfying $\sigma_{\boldsymbol{s}}\leq\xi^{1/q}$. Then for a given $n\in{\mathbb{N}}$, we selected a number $\xi_{n}$ such that the number of quadrature points in the grid $\Gamma(\Lambda_{\rm ev}(\xi_{n}))$ does not exceed $n$. Hence, due to the evenness of the indices in the set $\Lambda_{\rm ev}(\xi_{n})$ we obtained the improved rate 2/q1/22/q-1/2 and that the sparsity of $\Lambda_{\rm ev}(\xi_{n})$ is much higher then that of $\Lambda_{m(n)}$.

### v15

@@ 4 Integration
$\begin{split}\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}v\right\|_{X^{1}}\ &\leq\ \left\|v_{\rm ev}-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}\right\|_{{\mathcal{L}}_{1}(X^{1})}\\ \ &\leq\ C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2\end{cases}.\end{split}$
@@
$G_{2}(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq\xi\big\}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>1/q_{2}-1/2,\end{cases}$ (4.9)
@@ 28    $k\leftarrow k+1$
For $\xi>1$, let the set $G_{\rm ev}^{*}(\xi)$ be defined by
$G_{\rm ev}^{*}(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{\rm ev}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}/2}\leq\xi\big\}\ &{\rm if}\ \alpha\leq 2/q_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{\rm ev}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}/2}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big\}\ &{\rm if}\ \alpha>2/q_{2}-1/2,\end{cases}$ (5.6)
for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$, where
$\tau:=\frac{4\alpha}{4-q_{2}},\qquad\vartheta:=\frac{4}{4-q_{2}}\left(\frac{2}{q_{1}}-\frac{1}{2}\right).$ (5.7)
@@
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. Let the set $G_{\rm ev}^{*}(\xi)$ be defined as in (5.6). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}^{*}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}v\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{2}-1/2.\end{cases}$ (5.8)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}^{*}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{2}-1/2.\end{cases}$ (5.9)
@@
Observe that the rate in (5.8) and (5.9) can be improved as $\min(\alpha,\frac{2}{q_{1}}\frac{\alpha}{\alpha+\delta})$ if the sequences $(\|u_{\boldsymbol{s}}\|_{V})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ and $(\|u_{\boldsymbol{s}}\|_{W^{r}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ have $\ell_{p_{1}}$- and $\ell_{p_{r}}$-summable majorant sequences, respectively, where $1/p_{1}=1/q_{1}+1/2$ and $1/p_{r}=1/q_{r}+1/2$. Similarly to ${\mathcal{I}}_{G(\xi_{n})}$, the quadrature operator ${\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}$ can be represented in the form of a multilevel integration method with $k_{n}$ levels:
${\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}\ =\ \sum_{k=0}^{k_{n}}\delta_{k}Q_{\Lambda_{{\rm ev};k}^{*}(\xi_{n})},$
@@
$\Lambda_{{\rm ev};k}^{*}(\xi):=\ \begin{cases}\big\{{\boldsymbol{s}}\in{\mathbb{F}}_{\rm ev}:\,\sigma_{2;{\boldsymbol{s}}}^{q_{2}/2}\leq 2^{-k}\xi\big\}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big\{{\boldsymbol{s}}\in{\mathbb{F}}_{\rm ev}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}/2}\leq\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>1/q_{2}-1/2.\end{cases}$
For $\xi>1$, let the set $\Lambda_{\rm ev}^{*}(\xi)$ be defined by
$\Lambda_{\rm ev}^{*}(\xi):=\{{\boldsymbol{s}}\in{\mathbb{F}}_{\rm ev}:\,\sigma_{\boldsymbol{s}}^{q/2}\leq\xi\}.$ (5.10)
@@
Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let the set $\Lambda_{\rm ev}^{*}(\xi)$ be defined by (5.10) for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}^{*}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}^{*}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (5.11)
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}^{*}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}^{*}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.12)
@@
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{2}^{*}(\xi)$ be the set defined by
$G_{2}^{*}(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}/2}\leq\xi\big\}\quad&{\rm if}\ \alpha\leq 2/q_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}/2}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>2/q_{2}-1/2,\end{cases}$ (5.13)
for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$, where $\tau$ and $\vartheta$ are as in (5.7).
Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{2}^{*}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}v({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}^{*}(\xi_{n})}v\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{2}-1/2.\end{cases}$ (5.14)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{2}^{*}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}^{*}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{2}-1/2.\end{cases}$ (5.15)
@@
Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $\Lambda_{2}^{*}(\xi)$ be the set defined by
$\Lambda_{2}^{*}(\xi):=\{{\boldsymbol{s}}\in{\mathbb{F}}_{2}:\,\sigma_{\boldsymbol{s}}^{q/2}\leq\xi\}$
for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}^{*}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{R}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}^{*}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (5.16)
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}^{*}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}^{*}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.17)
@@
(i) As noticed in Section 4, the sparsity of the grids $\Gamma(\Lambda_{{\rm ev};0}^{*}(\xi))$ and $\Gamma(\Lambda_{\rm ev}^{*}(\xi))$ of the evaluation points in the quadrature operators ${\mathcal{Q}}_{G_{\rm ev}^{*}(\xi)}$ and $Q_{\Lambda_{\rm ev}^{*}(\xi)}$ are much higher than the sparsity of the grids $\Gamma(\Lambda_{0}(\xi))$ and $\Gamma(\Lambda(\xi))$ of the evaluation points in the generating interpolation operators ${\mathcal{I}}_{G(\xi)}$ and $I_{\Lambda(\xi)}$.
@@
(iii) Since the use and analysis of non-adaptive construction methods for sparse-grid interpolation are important, let us compare in details our methods in Corollary 5.2 with those which has been also discussed in [4]. To construct a quadrature of the form $Q_{\Lambda}$, the author of the last work used the set $\Lambda_{m}\subset{\mathbb{F}}$ of all indices $\boldsymbol{s}$} (including non-even) corresponding to the $m$ smallest values of $\sigma_{\boldsymbol{s}}$. The number $n=n(m)$ of quadrature points in $\Lambda_{m}\subset{\mathbb{F}}$ is estimated as $n\leq Cm^{2}$ [19, Proposition 3.16]. This lead to the rate $\frac{1}{2}(1/q-1/2)$. In the present paper, we used the set $\Lambda_{\rm ev}^{*}(\xi)\subset{\mathbb{F}}_{\rm ev}$ of all only even indices $\boldsymbol{s}$} by thresholding $\sigma_{\boldsymbol{s}}\leq\xi^{1/q}$. Formally, this is similar to choosing all even indices $\boldsymbol{s}$} corresponding to the smallest values of $\sigma_{\boldsymbol{s}}$ satisfying $\sigma_{\boldsymbol{s}}\leq\xi^{1/q}$. Then for a given $n\in{\mathbb{N}}$, we selected a number $\xi_{n}$ such that the number of quadrature points in the grid $\Gamma(\Lambda_{\rm ev}^{*}(\xi_{n}))$ does not exceed $n$. Hence, due to the evenness of the indices in the set $\Lambda_{\rm ev}^{*}(\xi_{n})$ we obtained the improved rate 2/q1/22/q-1/2 and that the sparsity of $\Lambda_{\rm ev}^{*}(\xi_{n})$ is much higher then that of $\Lambda_{m(n)}$.